The perimeter of a rectangular backyard is 6x + 6 yards. If the width is x yards, find a binomial that represents the length.
Answer by stanbon(75887)

The perimeter of a rectangular backyard is 6x + 6 yards. If the width is x yards, find a binomial that represents the length.
---------
Perimeter = 2(length + width)
6x+6 = 2(L + x)
3x+3 = L + x
2x+3 = Length
==================
Cheers,
Stan H.